OpenStudy (anonymous):

Slope = - 2, passing through (0, -3)

OpenStudy (texaschic101):

In y = mx + b form, the m stands for the slope and you can use your points (x,y) and sub...and since we have the slope, we need to find the y int, which is b y = mx + b slope(m) = -2 (0,-3)...x = 0 and y = -3 now we sub -3 = -2(0) + b -3 = 0 + b -3 = b so your equation is : y = -2x - 3 in slope intercept form

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

No thank you